OpenStudy (anonymous):

How to calculate ∜(i+1) ?

OpenStudy (ivancsc1996):

I don't think you can do it with algebra. I guess a calculator.

OpenStudy (amistre64):

\[(1+i)^{1/4}\] might help to do some arg and mod\[mod~(\cos(arg)+i~\sin(arg))\] given a+bi mod=sqrt{a^2+b^2} arg = inverse tan(b/a)

OpenStudy (amistre64):

then its just a matter of mod^(1/4), times [ cos(4arg) + i sin(4 arg)] if memory serves

OpenStudy (ivancsc1996):

Wow, what was that.

OpenStudy (anonymous):

Well, that's something too complicated! There should be an easier way, because it's an exam task.

OpenStudy (amistre64):



OpenStudy (amistre64):

there are different ways to approach it ...

mathslover (mathslover):

Using polar form may be easy here?

mathslover (mathslover):

\(\sqrt{1+i} ~\textbf{in polar form is : } ~ \sqrt{2} e^{\cfrac{i\pi}{4}} \)

mathslover (mathslover):

Related discussion here : http://openstudy.com/study#/updates/4e353f760b8ba7b2da426553 The above post has some mistake, it is "1+i" and not "\(\sqrt{1+i} \) "